How many atoms of chlorine are in a 30.0gmg tablet of Pseudoephedrine Hydrochloride?
1
Expert's answer
2018-02-13T06:28:00-0500
Chemical formula of Pseudoephedrine Hydrochloride is C 10 H 16 ClNO.
M (C 10 H 16 ClNO) = 201.7 g/mol
1 mole contains 6,02·10 ^23 of atoms.
n (C 10 H 16 ClNO) = 0.03 / 201.7 = 0.00015 moles
Chlorine takes 1/29 of all atoms in this substance. Therefore:
Number of Chlorine = 0.00015/29 · 6,02·10 ^23 = 8,96 · 10^ 19